An internet service provider states that the average number of hours its customers are online each day is 4.01. Suppose a random sample of 14 of the companys customers is selected and the average number of hours that they are online each day is measured. The sample results are shown in the accompanying table.
_____________________
3.85. 3.02. 6.16. 2.59. 3.71. 4.26. 8.07
3.63. 6.42. 1.37. 6.38. 7.89. 7.31. 1.19
_____________________
Based on the sampling of 14 customers, how much sampling error exists?
Answer #1

sample standard deviation

x̅ = 4.7

From the given data we can calculate sample standard deviation s = 2.32

Sampling error = s/√n = 2.32/√14 = 0.62